OpenStudy (anonymous):

How many atoms are there in 8.3 g Si?

OpenStudy (anonymous):

Do you mean the number of moles?

OpenStudy (anonymous):

No, I think that you have to convert from grams to moles to particles (atoms, molecules) I just don't remember whether to multiply or divide

OpenStudy (anonymous):

well since g was given, in order to find the no. of moles, you need to find the molar mass of si ( just type it in on google). The formula is \[\frac{ mass}{ molar mass }\] and from there you can find the no. of molecules given \[\frac{ no. of moles }{ avogadros }\] by the way avogadros number is 6.02 \[\times 6.02\times 10^{23}\] that's about it!

OpenStudy (anonymous):

correction its one 6.02 forget about the other one

OpenStudy (anonymous):

Did this help?

OpenStudy (anonymous):

When you're referring to mass/molarmass, the numerator would be the grams (8.3)? There fore the equation would be \[\frac{ 8.3 }{ 28.0855 }\] ?

OpenStudy (anonymous):

yes correct!

OpenStudy (anonymous):

Alright, thanks so much for everything! :D

OpenStudy (anonymous):

your welcome! :)